OpenStudy (anonymous):

(2n to the 4th power h to the 3rd power) 4 can someone please help me out here!!!!!!!!!!

OpenStudy (kc_kennylau):

\[\Large(2n^4h^3)^4\] ASCII notation: (2n^4h^3)^4

OpenStudy (kc_kennylau):

\[\Large(2n^4h^3)^4=2^4(n^4)^4(h^3)^4=16n^{16}h^{12}\]

OpenStudy (anonymous):

thank you!

OpenStudy (kc_kennylau):

no problem :)

OpenStudy (anonymous):

(-5p to the fourth power z to the sixth power u) 3 hint: apply the 3 exponent to each part..... I am lost please help!!!

OpenStudy (kc_kennylau):

\[\Large(abc)^n=a^nb^nc^n\]